A stone, thrown up is caught by the thrower after 6s. How high did it go and where was it 4 s after start? g = 9.8 m/s2

Grade:Select Grade

1 Answers

noogler
8 years ago
use formula t=2u/g.............find u=initial velocity
now use for h=u2/2g...............to find its height
to calculate height at t=4................... use formula hl=ut+1/2gt2